solve z^2 = -z+1 by using the quadratic formula
Answer by flame8855(424)

z^2 +z-1=0
a = 1 b =1 c = -1
delta = b^2-4ac = 1-4(1)(-1)= 5
x1= (-1-sqrt5)/2
x2= (-1+sqrt5)/2